# Real Noncommutative Convexity II: Extremality and nc convex functions

### v2

@@ 5. Convex multivalued functions and convex envelopes
6. Appendix by Travis Russell on nc extreme points
The following is based on the proofs from the unpublished note [3] without significant change, and is included for the readers convenience mainly so as to see that there are no hidden real vs. complex issues.
If $\varphi:\mathcal{S}\to B(H)$ and $\psi:\mathcal{S}\to B(K)$ are ucp maps on a real operator system, then $\varphi\leq\psi$ below means that $H\subseteq K$ and $\varphi$ is the restriction of $\psi$ to $H$ (i.e. $\psi$ is a dilation of the map $\varphi$). As usual $\varphi$ is maximal if every such $\psi$ is a trivial dilation as defined in Section 2. Let $\varphi:\mathcal{S}\to B(H)$ be a ucp map, and let $F\subset S\times H$. We say that $\varphi$ is maximal on $F$ if whenever $\psi\geq\varphi$, then $\psi(x)h=\varphi(x)h$ for all $(x,h)\in F$. If $\varphi:\mathcal{S}\to B(H)$ is maximal on $S\times H$, then it is maximal as defined above.
Theorem 6.1.
Every real ucp map $\varphi:\mathcal{S}\to B(H)$ can be dilated to a maximal ucp map $\widetilde{\varphi}:\mathcal{S}\to B(K)$.
Proof.
We sketch the proof that the arguments in the unpublished [3] (Lemma 1.2 and Theorem 1.3) apply for real operator systems. We make use of the direct (inductive) limit of Hilbert spaces repeatedly. Suppose $H_{1}\subseteq H_{2}\subseteq\dots$ is a sequence of real Hilbert spaces. Define an inner product on the disjoint union $\hat{H}=\bigsqcup H_{n}$ as follows: if $h\in H_{n}$ and $k\in H_{m}$, define $\langle h,k\rangle$ to be the inner product in $H_{k}$ where $k\geq n,m$. This is well defined since if $k\geq n,m$ an $l\geq n,m$, either $k\leq l$ or $l\leq k$, and the inner product in the smaller space agrees with the inner product in the larger space (since either $H_{l}\subseteq H_{k}$ or $H_{k}\subseteq H_{l}$). The vector space structure on $\hat{H}$ is defined similarly by taking (finite) linear combinations of vectors in a suitably large $H_{n}$. That the inner product is bilinear and positive definite on $\hat{H}$ follows, since each $H_{n}$ is a Hilbert space. It follows that the completion $H_{\infty}$ is a real Hilbert space. If $S$ is a real operator system and $\phi_{n}:S\to B(H_{n})$ are ucp with $\phi_{1}\leq\phi_{2}\leq\dots$, then for every $x\in S$ and $h_{n},k_{n}\in H_{n}$, then $\langle\phi_{\infty}(x)h_{n},k_{n}\rangle:=\langle\phi_{n}(x)h_{n},k_{n}\rangle$ defines a bounded bilinear form on $H_{\infty}$ so that $\phi_{\infty}(x)\in B(H_{\infty})$ is well-defined. It is easily verified that $\phi_{\infty}:S\to B(H_{\infty})$ is ucp with $\phi_{\infty}\geq\phi_{n}$ for every $n$.
We next show that for each $(x,h)\in S\times H$, there exists a dilation $\psi\geq\varphi$ maximal on $\{(x,h)\}$. As $\|\psi(x)h\|\leq\|x\|\|h\|$ for all $\psi\geq\varphi$, we have $C=\sup\{\|\psi(x)h\|:\psi\geq\varphi\}<\infty$. Inductively choose $\varphi\leq\varphi_{1}\leq\varphi_{2}\dots$ such that $\|\varphi_{n}(x)h\|\geq C-(1/n)$. Let $\varphi_{\infty}:\mathcal{S}\to B(H_{\infty})$ be the unique dilation to $H_{\infty}$. Then if $\psi\geq\varphi_{\infty}$,
$\|\psi(x)h\|\geq\|\varphi_{\infty}(x)h\|\geq\|\varphi_{n}(x)h\|\geq C-(1/n)\geq\|\psi(x)h\|-1/n$
for all $n$. Hence $\|\psi(x)h\|=\|\varphi_{\infty}(x)h\|$. Assume $\psi:\mathcal{S}\to B(K)$ with $H_{\infty}\subseteq K$. Since $P_{\infty}\psi(x)h=\varphi_{\infty}(x)h$, by the pythagorean identity
$\|\psi(x)h-\varphi_{\infty}(x)h\|^{2}=\|(I-P_{\infty})\psi(x)h\|^{2}=\|\psi(x)h\|^{2}-\|\varphi_{\infty}(x)h\|^{2}=0.$
Therefore $\varphi_{\infty}$ is maximal on $(x,h)$.
It then follows by transfinite induction that there exists $\psi:\mathcal{S}\to B(K)$, $\psi\geq\varphi$, which is maximal on $S\times H$. To see this, choose a well-ordering of $S\times H$. There exists a dilation $\phi_{0}:\mathcal{S}\to B(H_{0})$ of $\varphi$ which is maximal on the least element $(x_{0},h_{0})$, by the previous paragraph. Let $\lambda=(x_{\alpha},h_{\alpha})\in S\times H$. If $\lambda$ has a predecessor $(x_{\alpha-1},h_{\alpha-1})$, then apply the previous paragraph to get $\phi_{\alpha}\geq\phi_{\alpha-1}$ maximal on $(x_{\alpha},h_{\alpha})$. If $\lambda$ is a limit ordinal with $\alpha_{1}\leq\alpha_{2}\leq\dots,$ $\alpha_{n}\to\alpha$, and $\phi_{\alpha_{n}}:\mathcal{S}\to B(H_{\alpha_{n}})$ the corresponding dilations, then $\phi_{\infty}:\mathcal{S}\to B(H_{\infty})$ satisfies $\phi_{\infty}\geq\phi_{\alpha_{n}}$ for each $n$. By the previous paragraph again, $\phi_{\infty}$ has a dilation $\phi_{\alpha}\geq\phi_{\infty}$ maximal on $(x_{\alpha},h_{\alpha})$. By transfinite induction, we obtain a dilation $\psi:\mathcal{S}\to B(K)$ maximal on $S\times H$.
To finish, we may use the previous paragraph to construct a sequence of dilations $\varphi_{n}:\mathcal{S}\to B(H_{n})$ with $H\subseteq H_{1}\subseteq H_{2}\subseteq\dots$ such that $\varphi\leq\varphi_{1}\leq\varphi_{2}\leq\dots$ and such that $\varphi_{n}$ is maximal on $\mathcal{S}\times H_{n-1}$ (with $H=:H_{0}$). We claim $\varphi_{\infty}:\mathcal{S}\to B(H_{\infty})$ is maximal on $S\times H_{\infty}$ and hence is maximal. Clearly $\varphi_{\infty}$ is maximal on $S\times H_{n}$ for every $n$. Thus it is maximal on $S\times\bigcup H_{n}$. If $\{h_{k}\}\subseteq\bigcup H_{n}$ and $h_{k}\to h$ in $H_{\infty}$, then for any $\psi\geq\varphi_{\infty}$, $\psi(x)h=\lim\psi(x)h_{n}=\lim\varphi_{\infty}(x)h_{n}=\varphi_{\infty}(x)h$. So $\varphi_{\infty}$ is the desired maximal dilation. ∎
@@ 6. Appendix by Travis Russell on nc extreme points
More detailed proof of Theorem 2.20. As said (together with a few more specifics) below Theorem 2.20, we will use below that various properties of irreducible representations and corresponding GNS representations from [17], hold in the real case. We amplify Davidson and Kennedy’s argument, in part to see that there are no hidden real vs. complex issues. We will also use the following modified version of [14, Proposition 2.2.9]: Let $K$ be a closed nc-convex set over a dual operator space $E$. Suppose there is a net of contractions $\{\alpha_{i}\in M_{n,n_{i}}\}$ and a net $\{x_{i}\in K_{n_{i}}\}$ such that $\lim\alpha_{i}\alpha_{i}^{*}=I_{n}$ and such that $\lim\alpha_{i}x_{i}\alpha_{i}^{*}=x\in M_{n}(E)$. Then $x\in K_{n}$. To see this, let $\alpha_{i}=U_{i}P_{i}$ be a polar decomposition for $\alpha_{i}$, where $U_{i}\in M_{n,n_{i}}$ is a partial isometry and $P_{i}\in M_{n_{i}}$ is a positive operator. Let $\beta_{i}=U_{i}(I_{n_{i}}-P_{i}^{2})^{1/2}\in M_{n,n_{i}}$. Then $\alpha_{i}\alpha_{i}^{*}+\beta_{i}\beta_{i}^{*}=U_{i}U_{i}^{*}$, which is a projection in $M_{n}$. Let $m_{i}\leq n$ be a cardinal and $V_{i}\in M_{n,m_{i}}$ be a partial isometry such that $V_{i}V_{i}^{*}+U_{i}U_{i}^{*}=I_{n}$. Then because $\alpha_{i}\alpha_{i}^{*}\to I_{n}$ in $M_{n}(E)$, $\beta_{i}\beta_{i}^{*}\to 0$ and $V_{i}V_{i}^{*}\to 0$ in $M_{n}(E)$. Fix $y\in K_{1}$. Since $K$ is closed under nc-convex sets, $\alpha_{i}x_{i}\alpha_{i}^{*}+\beta_{i}(y\otimes I_{n_{i}})\beta_{i}^{*}+V_{i}(y\otimes I_{m_{i}})V_{i}^{*}\in K_{n}$ for every $i$. Since

### v3

@@ 5. Convex multivalued functions and convex envelopes
6. Appendix by Travis Russell: A more detailed proof of Theorem 2.20.
@@
As said (together with a few more specifics) below Theorem 2.20, we will use below that various properties of irreducible representations and corresponding GNS representations from [19], hold in the real case. We amplify Davidson and Kennedy’s argument, in part to see that there are no hidden real vs. complex issues. We will also use the following modified version of [16, Proposition 2.2.9]: Let $K$ be a closed nc-convex set over a dual operator space $E$. Suppose there is a net of contractions $\{\alpha_{i}\in M_{n,n_{i}}\}$ and a net $\{x_{i}\in K_{n_{i}}\}$ such that $\lim\alpha_{i}\alpha_{i}^{*}=I_{n}$ and such that $\lim\alpha_{i}x_{i}\alpha_{i}^{*}=x\in M_{n}(E)$. Then $x\in K_{n}$. To see this, let $\alpha_{i}=U_{i}P_{i}$ be a polar decomposition for $\alpha_{i}$, where $U_{i}\in M_{n,n_{i}}$ is a partial isometry and $P_{i}\in M_{n_{i}}$ is a positive operator. Let $\beta_{i}=U_{i}(I_{n_{i}}-P_{i}^{2})^{1/2}\in M_{n,n_{i}}$. Then $\alpha_{i}\alpha_{i}^{*}+\beta_{i}\beta_{i}^{*}=U_{i}U_{i}^{*}$, which is a projection in $M_{n}$. Let $m_{i}\leq n$ be a cardinal and $V_{i}\in M_{n,m_{i}}$ be a partial isometry such that $V_{i}V_{i}^{*}+U_{i}U_{i}^{*}=I_{n}$. Then because $\alpha_{i}\alpha_{i}^{*}\to I_{n}$ in $M_{n}$, we have $\beta_{i}\beta_{i}^{*}\to 0$ and $V_{i}V_{i}^{*}\to 0$ in $M_{n}$. Fix $y\in K_{1}$. Since $K$ is closed under nc-convex combinations, $\alpha_{i}x_{i}\alpha_{i}^{*}+\beta_{i}(y\otimes I_{n_{i}})\beta_{i}^{*}+V_{i}(y\otimes I_{m_{i}})V_{i}^{*}\in K_{n}$ for every $i$. Since e.g. $\beta_{i}\beta^{*}_{i}\to 0$ SOT, we have